The displacement x of a body of mass 1 kg on a horizontal smooth surface has the function of the time t is given by x=t⁴/4. The work done in first one second is

Grade:11

1 Answers

Arun
5 years ago
x = t^4 / 4  meters               m = 1 kg
dx = t²^3dt
dx/dt = v = t^3
d²x/dt² = a = 3 t^2

force acting on the body = m a = 3 t^2  Newtons
 
now fornwork done integrate F. dx